# ED4: Explicit Data-level Debiasing for Deepfake Detection

### v1

@@ Introduction
Table 5: Results (AUC) of adding ED4 to existing deepfake detection methods for plug-and-play effectiveness.
@@
Plug-and-Play with Previous Deepfake Detectors
@@
Considering our debiasing method is explicitly applied at the data level, it can be directly implemented to existing SOTA methods for deepfake detection. Therefore, we reproduce EfficientNetB4 (Effnb4) (Tan and Le 2019), SPSL (Liu et al. 2021), and Capsule (Nguyen, Yamagishi, and Echizen 2019) and then adding ED4 to them. Tab. 5 shows that with the additive implementation of ED4, all detectors exhibit significant enhancements in performance. This substantiates that our method is plug-and-play that can be conveniently applied to other methods to enhance their effectiveness.
Figure 4: Comparisons of CAM for different methods.
Qualitative Results
@@
In this section, we visualize the Class Activation Map (CAM) via Grad-CAM++ (Chattopadhay et al. 2018) to discuss the focal regions of different detectors. UCF (Yan et al. 2023a) is the most advanced method that addresses model bias by feature disentanglement via implicit network design. In Fig. 4, we provide the CAM of images with four different conditions. Specifically, Spatial Shuffle (Fig. 4 (a)) and ClockMix (Fig. 4 (b)) are operations introduced by our method. Local Forgery (Fig. 4 (c)) is achieved by replacing eyes in the real faces with the fake ones following (Chen et al. 2022). Images in Cross-Dataset (Fig. 4 (d)) are from CDFv1 while detectors are trained on FF++.
@@
Suffering from spatial bias, the Baseline is inertly focusing on the relatively central regions despite the facial regions being relocated or locally replaced. UCF can perceive forgery clues in a wilder region of the image center (the 2nd rows in Fig. 4 (a) and (c)), but it also struggles to precisely locate the forgery artifacts (Fig. 4 (b) and the 1st rows in Fig. 4 (a) and (c)). Conversely, our method can adaptively focus on the local forgery artifacts at diverse image locations. Fig. 4 (d) indicates the challenges for the existing methods to detect the common artifacts in cross-dataset settings. Namely, the Baseline fails to perceive forgery, while UCF may overlook the actual forgery regions (the 2nd row in Fig. 4 (d)). Conversely, our method exhibits superior sensitivity to the cross-dataset common forgery artifacts.
Conclusion
@@
Supplementary Materials
More Implementation Details
Detailed Experimental Setting
Here, we provide the implementation configuration we applied for the experiments of the main paper. To ensure fair comparisons, all methods present in our paper follow the same configurations from DeepfakeBench (Yan et al. 2023b) as we provided below. Dlib (dlib) is used for face extraction and alignment with a cropping margin of 15%. 32 frames are extracted from each video and all frames are cropped to the size of 256 $\times$ 256. For training, traditional data augmentations are deployed including rotating, Gaussian noise, saturation adjusting, and quality adjusting.
Detailed Implementations of the Adversarial Generator
Due to length limitations, we did not provide a detailed introduction to the proposed adversarial generator in the main manuscript. Here, we comprehensively discuss the architectural details of the adversarial generator and the rationale for applying linear programming to the probability metric $m$ via the Hungarian algorithm. Firstly, our adversarial generator is based on an Xception backbone. Given a granularity $g$, the number of divided patches is $N=g^{2}$, and hence the shape of the output possibility metric $\mathbf{m}$ is $N\times N$, where the first $N$ refers to the number of patches in the shuffled image, the second one refers to the possible shuffled location of patches. Such an effect can be easily achieved by deploying a linear layer that outputs $N\times N$ after the Xception backbone and then reshaping the output.
However, directly using $\mathbf{m}$ to guide the shuffling of images could be problematic. Specifically, each region should be assigned to a unique patch, that is, the predicted region index for each patch must be distinct. In fact, this mutual exclusivity of indices aligns with solving polynomials, except we are maximizing polynomials instead of minimizing them, which could be easily transformed by adding a minus sign to $\mathbf{m}$. Therefore, the problem can be formulated as:
where $x_{i,j}\in\{0,1\}$ and
Figure 5: Visualized example of transforming generator output to shuffle permutation.
Robustness Evaluation
Figure 6: Robustness on Unseen Perturbations
Robustness Against Unseen Perturbations
As shown in Fig. 7, we introduce block-wise and blur for the robustness evaluation against unseen perturbations. To comprehensively investigate the robustness, we set the intensity of perturbations to gradually increase. Obviously, our method exhibits superior robustness due to the improved reception of forgery clues.
Figure 7: Robustness against spatial deviations. "Variations" indicates the variation of prediction scores when applying different spatial transformations. The dotted lines represent the average variation
Table 6: Comparison with more augmentation methods for spatial bias. Local represents the introduction of local forgery. Region denotes introducing region variations. Info-I denotes Informative pixel Integrity.
Method Local Region Info-I DFD CDFv2 DFDC Avg.
w/o AdvSCM (Baseline) 0.8200 0.7712 0.7296 0.7736
GridMask 0.8408 0.7838 0.7312 0.7853
AttrMask 0.8005 0.7139 0.6759 0.7301
RandomCrop 0.8415 0.8150 0.7235 0.7933
Shuffle 0.8765 0.8394 0.7427 0.8195
Robustness Against Spatial Deviations
We provide a prediction stability study to investigate the robustness against spatial deviation. Specifically, we quantify the variation in prediction confidence for each image after applying the spatial deviation and subsequently depict these variations as a histogram in Fig. 7. Mathematically, the deviation value at each image can be written as $|\mathcal{N}(\mathbf{I})-\mathcal{N}(\mathcal{T}_{s}(\mathbf{I}))|$, where $\mathcal{N}(\cdot)$ is the detection network, $\mathbf{I}$ is the original image, and $\mathcal{T}_{s}$ is one of the spatial transforming operation. A larger variation indicates a worse detection stability. For the baseline detector, a considerable number of detection results suffer from the spatial deviation. Especially for Crop, over one-third of the variations exceed 0.5, indicating a notable number of detection results are reversed to the opposite after spatial deviation. In contrast, with the help of AdvSCM, our detector can maintain strong robustness against different spatial deviations.
Table 7: Investigation for mixing images with different labels. Avg. denotes the Average performance in the cross-dataset evaluation.
Method DFD CDFv2 DFDC Avg.
Baseline 0.7994 0.7454 0.7089 0.7512
w/o Real+Real 0.7177 0.6959 0.6361 0.6832
w/o Real+Fake 0.8481 0.8045 0.7255 0.7927
w/o Fake+Fake 0.8560 0.7931 0.7247 0.7913
Ours 0.8765 0.8394 0.7427 0.8195
More Ablation Study
Comparison with More Augmentation Methods for Spatial Bias
As shown in Tab. 6, we introduce three approaches to replace Shuffle that may contribute to spatial bias removal: 1) GridMask (gridmask), which is a widely-used augmentation that randomly masks a certain ratio of images. 2) AttrMask, which can mask out certain facial attributes guided by the facial landmarks. 3) RandomCrop, which alters the relative location of faces in the images. The w/o AdvSCM can be treated as the baseline for which no spatial transformation is deployed. The masking-based methods can break the structural forgery into local ones, but cannot improve the model sensitivity to artifacts at arbitrary regions. While RandomCrop can alter the forgery location to some extent but still maintain the forgery structure. Meanwhile, since masking and cropping lead to the reduction of informative pixels, the performances achieved by these methods are not as pronounced as our S-Shuffle. Especially for AttrMask, despite masking certain facial attributes that may enhance the model sensitivity to local forgery, it significantly reduces the pixels that likely contain forgery information and thus leads to performance degradation. Moreover, this information reduction further leads to inherent inconsistency in extracted features, thereby rendering the employment of Consistency Loss inappropriate. In contrast, deploying Shuffle can uniformly address spatial bias while maintaining informative pixel integrity. Therefore, we use Shuffle as the basic spatial operation to construct the proposed adversarial spatial consistency module.
Table 8: Cross-dataset Evaluations on more various datasets. The metric is frame-level AUC and all models are trained on FF++ (Rossler et al. 2019).
Methods WDF FAVC DiffSwap E4S BlendFace UniFace DFL Avg.
EfficientNet 0.7275 0.8404 0.7956 0.6514 0.7813 0.7775 0.7347 0.7583
Xception 0.7705 0.8681 0.8110 0.6554 0.7911 0.8282 0.7199 0.7777
SPSL 0.7023 0.7689 0.7784 0.5987 0.6936 0.6915 0.6548 0.6983
UCF 0.7738 0.8631 0.8511 0.6916 0.7850 0.7867 0.7433 0.7849
Ours 0.7682 0.8879 0.8781 0.7201 0.8511 0.8280 0.7794 0.8161
The Effectiveness of Conducting ClockMix on Arbitrary Images
To investigate the influence of conducting ClockMix on arbitrary images, we perform an ablation study with three ablation variants, that is, w/o Real+Real. w/o Real+Fake, and w/o Fake+Fake. As shown in Tab. 7, w/o Real+Real will undermine the performance by indulging the detector to take the seaming patterns as fake. Such a result is similar to the “mixture of real images with the fake label (MRF)” that we present in the main paper except the degradation is slighter. Overall, the lack of any label combination in ClockMix can lead to considerable performance degradation, which indicates the superiority of mixing arbitrary images over existing label-specific and image-specific synthesis methods.
More Cross-dataset Evaluations
In this section, we provide results of our method on seven more datasets, including WildDeepfake (wilddeepfake) (WDF), FakeAVCeleb (fakeavceleb) (FAVC), {E4S (e4s), BlendFace (blendface), UniFace (uniface), DeepfaceLab (deepfacelab) (DFL)} from DF40 (df40), and {DiffSwap (diffswap)} from DiffusionFace (diffusionface). Considering most baseline methods from DeepfakeBench (Yan et al. 2023b) have not been reproduced on these advanced datasets, we carefully reproduce EfficientNet (Tan and Le 2019), Xception (Chollet 2017), SPSL (Liu et al. 2021), and UCF (Yan et al. 2023a) and evaluate them on these datasets for comparison. As shown in Tab. 8, it can be observed that our methods can be generalized to various datasets with superior effectiveness.

### v2

@@ V-E Robustness Evaluation
TABLE VIII: Results (AUC) of adding ED4 to existing deepfake detection methods for plug-and-play effectiveness.
@@
Figure 9: Class Activation Mapping (CAM)-based attention heatmaps for different methods on representative forgery images. ED4 produces more concentrated and accurate attention on manipulated regions across diverse challenging scenarios, providing visual evidence of its superior sensitivity and localization precision.
V-F Plug-and-Play with Previous Deepfake Detectors
@@
Considering our debiasing method is explicitly applied at the data level, it can be directly implemented to existing SOTA methods for deepfake detection. Therefore, we reproduce EfficientNetB4 (Effnb4) effnet , SPSL spsl , and Capsule capsule and then adding ED4 to them. Tab. VIII shows that with the additive implementation of ED4, all detectors exhibit significant enhancements in performance. This substantiates that our method is plug-and-play that can be conveniently applied to other methods to enhance their effectiveness.
V-G Qualitative Results
@@
In this section, we visualize the Class Activation Map (CAM) via Grad-CAM++ grad-cam++ to discuss the focal regions of different detectors. UCF ucf is the most advanced method that addresses model bias by feature disentanglement via implicit network design. In Fig. 9, we provide the CAM of images with four different conditions. Specifically, Spatial Shuffle (Fig. 9 (a)) and ClockMix (Fig. 9 (b)) are operations introduced by our method. Local Forgery (Fig. 9 (c)) is achieved by replacing eyes in the real faces with the fake ones following adv . Images in Cross-Dataset (Fig. 9 (d)) are from CDFv1 while detectors are trained on FF++.
@@
Suffering from spatial bias, the Baseline is inertly focusing on the relatively central regions despite the facial regions being relocated or locally replaced. UCF can perceive forgery clues in a wilder region of the image center (the 2nd rows in Fig. 9 (a) and (c)), but it also struggles to precisely locate the forgery artifacts (Fig. 9 (b) and the 1st rows in Fig. 9 (a) and (c)). Conversely, our method can adaptively focus on the local forgery artifacts at diverse image locations. Fig. 9 (d) indicates the challenges for the existing methods to detect the common artifacts in cross-dataset settings. Namely, the Baseline fails to perceive forgery, while UCF may overlook the actual forgery regions (the 2nd row in Fig. 9 (d)). Conversely, our method exhibits superior sensitivity to the cross-dataset common forgery artifacts. It should be noted that ED4 is not designed to explicitly localize the precise forgery regions, the capability shown in Fig. 9 actually emerges from the following characteristics of ED4:
  • The proposed ED4 introduces strong data-level augmentation that effectively disrupts spurious correlation-based shortcuts for learning. As a result, the model has to become more sensitive to forgery artifacts and is compelled, during training, to implicitly learn to distinguish forged regions in order to produce correct predictions.
  • During training, ED4 is not applied to all inputs indiscriminately; instead, it is introduced at a controlled ratio to enrich the diversity of training samples. Consequently, the model implicitly acquires the ability to more accurately attend to and localize the forged regions.
VI Limitations and Future Work
Limitation. ED4 is an explicit debiasing framework specifically designed for facial forgery detection. Accordingly, the proposed ClockMix requires pre-extracted facial landmarks to perform face alignment, ensuring a semantic consistency of facial regions during the mixing process. However, in broader AI-generated content (AIGC) detection scenarios, it is often infeasible to align semantically diverse images prior to applying ClockMix. As a result, ClockMix cannot be readily extended to general-purpose AIGC detectors and struggles to accommodate more diverse and unconstrained forgery scenarios.
Future Work. As semantic bias also poses a significant challenge in AIGC detection, the development of an explicit debiasing module tailored to this task remains an important direction for future research. One potential approach is to exploit the attention regions of the classifier to identify semantically salient areas within images. Semantic alignment could then be performed based on these regions, serving as a foundation for designing an explicit debiasing strategy adapted for AIGC detection.
VII Conclusion